2 Find the frequency of vibration for r= 3= cm, R=3 meters, and m= 10

any help will be appreciated.
Answer #1

for angle o ngrino T= mg sino (R-8) and I= Im (982) T= mg sino (R-1) = I do for small angle sino a mgo(R-8)= I do myle_v) er